### v1

# Distraction is All You Need for Multimodal Large Language Model Jailbreaking

Abstract
Multimodal Large Language Models (MLLMs) bridge the gap between visual and textual data, enabling a range of advanced applications. However, complex internal interactions among visual elements and their alignment with text can introduce vulnerabilities, which may be exploited to bypass safety mechanisms. To address this, we analyze the relationship between image content and task and find that the complexity of subimages, rather than their content, is key. Building on this insight, we propose the Distraction Hypothesis, followed by a novel framework called Contrasting Subimage Distraction Jailbreaking (CS-DJ), to achieve jailbreaking by disrupting MLLMs alignment through multi-level distraction strategies. CS-DJ consists of two components: structured distraction, achieved through query decomposition that induces a distributional shift by fragmenting harmful prompts into sub-queries, and visual-enhanced distraction, realized by constructing contrasting subimages to disrupt the interactions among visual elements within the model. This dual strategy disperses the model’s attention, reducing its ability to detect and mitigate harmful content. Extensive experiments across five representative scenarios and four popular closed-source MLLMs, including GPT-4o-mini, GPT-4o, GPT-4V, and Gemini-1.5-Flash, demonstrate that CS-DJ achieves average success rates of 52.40% for the attack success rate and 74.10% for the ensemble attack success rate. These results reveal the potential of distraction-based approaches to exploit and bypass MLLMs’ defenses, offering new insights for attack strategies.
Warning: This paper contains unfiltered content generated by MLLMs that may be offensive to readers.
Figure 1: Jailbreaking examples of Hades [10] and our proposed CS-DJ against GPT-4o. When using Hades, GPT-4o effectively rejects harmful questions. In contrast, CS-DJ successfully bypasses the model’s defense mechanisms by constructing multi-subimage visual inputs to distract the model, resulting in the acquisition of harmful responses. The harmful information is highlighted in red, while key prompts are emphasized in bold.
1 Introduction
Since the early 2020s, Multimodal Large Language Models (MLLMs) have emerged as a powerful tool for bridging natural language processing and computer vision [3], owing to their ability to integrate and process multimodal information efficiently [28].
Despite their impressive performance, one notable limitation of these methods is that the training materials may contain toxic content, potentially leading to inappropriate or harmful content [42, 4], the leakage of sensitive information [40], and even serious vulnerabilities in the model security and privacy protection. To this end, significant efforts have been devoted to using reinforcement learning from human feedback (RLHF) to align LLMs outputs, ensuring readiness for deployment in high-stakes domains [22, 36]. However, beyond Large Language Models (LLMs), the integration of visual inputs in MLLMs introduces a new challenge: securing models against vulnerabilities arising from newly integrated visual modalities, which remains a significant issue and lacks a comprehensive understanding [41, 18, 16]
In this study, rather than following the human feedback alignment research at the defense side, we focus on exploring an alternative approach, namely jailbreak attacks [37, 40, 19]. This line of research also seeks to provide insights into how safety concerns emerge, while, considering the attacker’s perspective. Specifically, it aims to actively uncover the defense vulnerabilities in current model by deliberately crafting inputs to manipulate MLLMs, leading to unintended and potentially harmful outputs [14]. Motivated by the fact that most current defense mechanisms rely on RLHF and are trained on human preference data [22], a natural hypothesis regarding the defense vulnerabilities of MLLMs is that the model may be susceptible to out-of-distribution (OOD) inputs.
Given that harmful textual content is already effectively detected in the field of LLMs, current jailbreak attack research on MLLMs primarily focuses on constructing OOD visual inputs, with two main approaches: image perturbation injection and prompt-to-image infection. The image perturbation injection involves subtly modifying images and combining them with query text to compose jailbreak prompts [38]. This attack leverages adversarial approaches [33] to introduce noisy OOD inputs that exploit the MLLM’s visual-text decision boundary [44]. Unfortunately, these methods typically require access to the gradient of the victim or surrogate model [35], which limits their applicability in closed-source models [32]. In contrast, prompt-to-image infection generates images with similar semantics by using malicious text. These generated images are then combined with query text to craft jailbreak prompts that compel the MLLM to produce harmful responses [6, 10]. Since these generated images differ in distribution from those used for MLLM safety alignment training, they may bypass security measures and achieve jailbreak [15, 29, 2]. Despite this, the extensive training of large MLLMs on vast datasets makes it challenging for image generation models to produce truly OOD images. A failure case is depicted in the upper portion of Figure 1.
Findings. Despite progress, it remains unclear how to fully leverage the visual space to construct OOD inputs for effective black-box MLLM jailbreaks. Building on the observation that most harmless alignment processes primarily involve simple images [6, 10, 23], we propose that increasing image complexity, potentially through the use of multiple subimages, could generate more effective OOD inputs for jailbreak attack. To verify this, we conduct several ablation analyses (§$\S$4.4), including varying the number of subimages and examining the relationship between image content and task. Our findings reveal that it is the complexity of the subimages, rather than their conceptual content, that drives the jailbreak success. Based on this insight, we propose the Distraction Hypothesis to explain the visual effects on jailbreak attacks against MLLMs.
Distraction Hypothesis. Encoding complex images in the input prompt increases token complexity/diversity, which raises the processing burden on MLLMs. This overload can weaken the model’s defenses, making it more prone to induce unintended outputs and improving jailbreak attack effectiveness.
Structure and contributions. Building on the Distraction Hypothesis, we further introduce a novel framework, namely Contrasting Subimage Distraction Jailbreaking (CS-DJ) (§$\S$3), which leverages this insight to strategically design OOD inputs and enhance the effectiveness of jailbreak attacks. Specifically, CS-DJ explores both textual and visual components of the input space, introducing two key methods: structured distraction via query decomposition (§$\S$3.1) and visual-enhanced distraction through contrasting subimages (§$\S$3.2). (i) The structured distraction component decomposes the original harmful query into multiple sub-queries, each representing different aspects or intermediate steps of the original query. This decomposition induces a distributional shift that disperses the model’s focus and weakens its ability to detect harmful content. (ii) The visual-enhanced distraction component manipulates the visual input by constructing multiple contrasting subimages, effectively disrupting interactions among visual elements during processing. Then, to guide the model’s response and enhance its susceptibility to the multi-level distractions, CS-DJ incorporates a carefully designed harmless prompt alongside the composite visual input, as illustrated in Figure 1. This integrated approach disrupts the model’s internal coherence and safety mechanisms, effectively bypassing its defenses.
To evaluate the effectiveness of CS-DJ, we conduct extensive experiments across five representative scenarios and four widely used closed-source MLLMs (§$\S$4). Our results demonstrate that CS-DJ outperforms state-of-the-art jailbreak attacks, achieving average success rates of 52.40% for attack success rate and 74.10% for ensemble attack success rate.
2 Related Work
2.1 Multimodal Large Language Models
In recent years, Multimodal Large Language Models (MLLMs) [34] have emerged as the leading solution for handling complex multimodal data, such as images[43] and audio [5]. Inspired by instruction-tuning algorithms in LLMs, many MLLMs are fine-tuned from pre-trained LLMs. Notable examples include LLaVA [13] and MiniGPT-4 [46], which utilize similar architectures by connecting pre-trained vision backbones to large language models via a linear layer. These models are then fine-tuned using ChatGPT-generated data, enabling deep integration between vision and language. To enhance performance and usability, LLaVA-1.5 [12] replaces the linear layer with MLPs and incorporates prompt-based response formats optimized for academic VQA datasets.
MM1 [20] and MoE-LLaVA [11] further introduce MoE [26] structure to achieve better performance, without increasing computational cost. However, the integration of emerging modalities further intensifies the challenges associated with the safety alignment.
To achieve this, MLLMs are typically fine-tuned using data annotated with human preferences to ensure that their outputs are helpful and harmless. Despite these efforts, this approach remains vulnerable to attacks using OOD inputs, which will be thoroughly explored in this work.
2.2 Jailbreak Attacks against MLLMs
Jailbreak attacks against MLLMs can be conducted through both visual and textual inputs, leading to complex and diverse attack patterns [8, 21]. Current jailbreak attack techniques can be broadly categorized into two types: image perturbation injection and prompt-to-image infection jailbreaks. Among them, image perturbation injection techniques are particularly effective for white-box jailbreak scenarios.
In particular, to bypass the model’s security detection mechanisms, Qi et al. [23] conceal harmful textual information within images by injecting invisible noise [45] through gradient optimization. Furthermore, JIP [25] conceals harmful image information within benign images through latent space semantic alignment. To enable jailbreaks on black-box MLLMs, Hades [10] investigates the influence of image harmfulness on model jailbreaks and proposes a prompt-to-image infection jailbreak attack method. It generates harmful images using diffusion models [27, 39, 19] and amplifies their effect through optimization. The generated image is then combined with the query text to form a jailbreak prompt [15, 17, 6], successfully achieving a jailbreak. In a word, existing MLLM jailbreak attack methods generally focus on crafting out-of-distribution (OOD) inputs. However, it remains unclear how to effectively leverage the visual space to construct OOD inputs for successful black-box MLLM jailbreaks. To address this, we propose CS-DJ, which employs distraction to guide the construction of OOD inputs.
3 Methodology
Figure 2: The framework of the proposed CS-DJ. Given a harmful query, CS-DJ employs a three-step process to execute a jailbreaking attack: (1) decompose the raw query into multiple sub-queries and transform them into images to introduce structural distraction, (2) retrieve contrasting images from a dataset as subimages to inject visual-enhanced distraction. (3) combine these images into a composite image as the visual input, aligned with a harmless instruction, to execute the jailbreaking attack.
Beyond the standard LLMs, MLLMs also encode visual content into token sequences, saying that the intrinsic interactions among visual elements can affect the model’s ability to defend against harmful prompts. However, most of the existing approaches [10, 15, 16] only take a single, complete image as input, resulting in notable similarities both (1) within image patches, and (2) between the image and the targeted jailbreak content. Such similarities may make the model likely to detect harmful inputs, potentially leading to jailbreak failure. To address these issues, we propose a novel jailbreak framework, called CS-DJ, as illustrated in Figure 2. CS-DJ leverages a multi-subimage structure to bypass the models’ intrinsic safety detection mechanism, with introducing multi-level distraction. Specifically, CS-DJ consists of two main components: structured distraction, achieved through query decomposition, and visual-enhanced distraction, realized through the construction of contrasting subimages. The resulting composite image input, combined with a carefully designed instruction, is fed into MLLMs to achieve jailbreak. Further details are described in the following sections.
3.1 Structured Distraction via Query Decomposition
MLLMs have strong defense ability against harmful textual inputs, reinforced by specific optimizations aligned with human preference data [9]. To circumvent these defenses, we introduce a structured distraction strategy that disrupts and fragments information alignment through query decomposition, thereby reducing the model’s overall sensitivity to harmful content. As illustrated on the left of Figure 2, we begin by decomposing a original harmful query $Q$ into multiple sub-queries, denoted as $\{Q_{s}^{(i)}\}_{i=1}^{m}$. Each sub-query addresses part of the original query from different aspects or intermediate steps.
$\{Q_{s}^{i}\}_{i=1}^{m}=\mathcal{G}(Q),$ (1)
where $\mathcal{G}$ is the auxiliary decomposition model and $m$ denotes the number of sub-queries. On one hand, this fragmented structure distracts the model’s focus, creating a distributional shift that obscures the original intent and disrupts its ability to detect harmful content, thereby bypassing its safety mechanisms. On the other hand, during the jailbreak execution phase, we carefully design prompts that guide the MLLM to address all sub-queries simultaneously. This strategy causes the model to divide its attention among multiple tasks rather than focusing on a single query. Further details are discussed in Section 3.3. In practice, we utilized Qwen2.5-3B-Instruct [31] for query decomposition.
The details of the query decomposition prompt are provided in the Appendix.
The next step involves converting the sub-queries into images:
${\{I_{s}^{i}\}}_{i=1}^{m}=\mathcal{T}(\{{Q_{s}^{i}\}}_{i=1}^{m}),$ (2)
where $\mathcal{T}$ denotes the transformation function that converts each sub-query $Q_{s}^{i}$ into a corresponding image $I_{s}^{i}$. This transformation introduces a new form of distraction by altering the input modality, complicating the model’s ability to link and identify harmful content patterns. Presenting textual prompts as visual elements obscures their original intent, further disrupting the model’s detection mechanisms.
3.2 Visual-Enhanced Distraction via Multi-subimage Construction
In contrast to prior work that emphasizes the harmfulness of a single image for MLLM jailbreaks, we focus on the impact of distraction. To enhance visual input distraction, we construct multiple contrasting visual elements, each treated as a subimage that collectively forms the input image. Since MLLMs represent visual content as token sequences similar to text, the distraction must include two aspects: (1) the distraction between the text query and the visual subimage, and (2) the distraction among the visual subimages themselves. Directly constructing an image with sufficient distraction relative to a text query is challenging. Therefore, we simplify this process by transforming it into an image retrieval problem, aimed at retrieving $k$ images from a dataset. These $k$ images are selected to have the lowest similarity to the query and to each other. Given the complexity of this optimization problem, we further approximate it for practical implementation. The procedure is illustrated in the middle of Figure 2. Specifically, we first encode the query $Q$ as a dense vector using the CLIP model:
$v(Q)=\text{CLIP}(Q).$ (3)
In practice, CLIP-ViT-L/14 [24] is utilized as the extractor. We then retrieve an image from a dataset that most contrasts with $v(Q)$, effectively by minimizing the cosine similarity as:
$I_{c}^{1}=\arg\min_{I\in\mathcal{D}}\textbf{cos}(v(Q),v(I)),$ (4)
where $\mathcal{D}$ denotes the image dataset, and cos($\cdot$,$\cdot$ , ⋅) represents the cosine similarity. Next, we proceed to retrieve the subsequent subimages:
$I_{c}^{j}=\arg\min_{I\in\mathcal{D}}\left(\textbf{cos}(v(Q),v(I))+\sum_{i=1}^{j-1}\textbf{cos}\left(v(I_{c}^{i}),v(I)\right)\right),$ (5)
where $j$ denotes the index of the current subimage being selected. This approach systematically ensures that each subimage has minimal similarity to both the query and the previously selected subimages, thereby maximizing contrast and enhancing the overall distraction effect for the MLLM jailbreak process.
3.3 Jailbreaking Execution
With the constructed set of subimages, we form a composite input that maximizes distraction for MLLMs. Specifically, the final composite image $I_{\text{comp}}$ is created by combining the subimages $\{I_{c}^{j}\}_{j=1}^{k}$ along with the transformed sub-queries $\{I_{s}^{i}\}_{i=1}^{m}$ :
$I_{\text{comp}}=\textbf{Combine}\left({\{I_{c}^{j}\}}_{j=1}^{k},{\{I_{s}^{i}\}}_{i=1}^{m}\right),$ (6)
where the function $\textbf{Combine}(\cdot,\cdot)$ represents the structured arrangement of the subimages and sub-query transformations.
This composite input $I_{\text{comp}}$, along with a harmless instruction $P$, is fed into a MLLM for processing:
$Y=\text{MLLM}(I_{\text{comp}},P),$ (7)
where $Y$ denotes the textual output of the MLLM.
The use of $I_{\text{comp}}$ with contrasting visual elements maximizes distraction, effectively disrupting the model’s internal alignment and coherence mechanisms. To complement $I_{\text{comp}}$, we carefully design $P$ to further enhance the distraction effect. As illustrated in Figure 1, $P$ comprises three main components: (1) a role-guiding section, (2) a task-guiding section, and (3) a visual-guiding section. The role-guiding strategy is commonly employed across various LLM tasks. The second component of $P$ directs the model to process multiple sub-queries simultaneously, increasing task complexity and dispersing attention across multiple tasks instead of focusing on a single one. Finally, $P$ includes a misleading instruction, informing the model that the visual subimages may contain useful information, further diverting its attention. By presenting these inputs together, the approach reduces the model’s ability to detect and mitigate harmful content, effectively bypassing its intrinsic safety mechanisms through multi-level distraction.
3.4 Distraction Distance Measurement
To quantify the effectiveness of our multi-subimage input for distraction, we introduce a metric called Distraction Distance. This metric evaluates interactions among the query $Q$ and contrasting visual images ${\{I_{c}^{j}\}}_{j=1}^{k}$, treating all as equally important nodes within a unified structure.
Each node is represented by its CLIP-encoded vector. For a given node, we calculate the $\text{L}_{2}$ distance between its vector and the vectors of all other nodes. The sum of these $\text{L}_{2}$ distances for each node is calculated across all other nodes to obtain the Distraction Distance:
$DL=\sum_{i=1}^{N}\sum_{j\neq i}||v_{i}-v_{j}||_{2},$ (8)
where $N$ denotes the total number of nodes, and $v_{i}$ represents the CLIP-encoded vector of node $i$. This approach accounts for both the number of nodes and their pairwise distances, providing a comprehensive measure of the dispersion and contrast within the multi-subimage structure. By maximizing this Distraction Distance, we amplify the overall distracting effect on MLLMs, thereby disrupting their ability to accurately detect harmful content.
4 Experiment
Victim Model Method Animal Financial Privacy Self-Harm Violence Average (%)
Attack Success Rate (ASR) $\uparrow$
GPT-4o-mini Hades 4.77 11.66 6.88 1.55 5.55 6.08
CS-DJ 26.55 80.00 77.11 34.77 70.55 57.80 (+51.72)
GPT-4o Hades 2.33 10.11 7.55 1.88 5.66 5.51
CS-DJ 15.33 70.77 58.55 15.77 50.77 42.24 (+36.73)
GPT-4V Hades 8.11 32.55 23.88 8.00 29.11 20.33
CS-DJ 11.66 76.22 59.77 22.22 57.33 45.44 (+25.11)
Gemini-1.5-Flash Hades 2.66 23.44 13.11 1.77 5.00 9.20
CS-DJ 22.55 88.88 79.11 47.00 83.00 64.11 (+54.91)
Ensemble Attack Success Rate (EASR) $\uparrow$
GPT-4o-mini Hades 10.66 21.33 13.33 4.66 10.00 12.00
CS-DJ 54.00 92.66 94.66 52.00 92.66 77.20 (+65.20)
GPT-4o Hades 5.33 16.00 12.00 6.00 11.33 10.13
CS-DJ 35.33 92.66 83.33 37.33 80.66 65.86 (+55.73)
GPT-4V Hades 18.66 63.33 54.00 24.66 64.66 45.06
CS-DJ 34.66 94.00 89.33 52.0 89.33 71.86 (+26.80)
Gemini-1.5-Flash Hades 4.66 34.00 20.00 4.66 10.66 14.79
CS-DJ 51.33 97.33 95.33 66.00 97.33 81.46 (+66.67)
Table 1: ASR and EASR results of CS-DJ and Hades on four closed-source MLLMs across different categories.
In this section, we first introduced our experimental settings which include datasets and evaluation metrics. Then, we carried out quantitative and qualitative comparisons between our method and state-of-the-art MLLM jailbreaking approaches. Additionally, ablation studies are performed to validate the distraction effect on MLLM jailbreaking from various perspectives.
4.1 Experiment Setup
Dataset. The datasets are primarily utilized to provide or construct visual inputs tailored to the requirements of different methods. For evaluating Hades, we followed the experimental setting outlined in its original paper, utilizing the Hades dataset [10] to supply the visual inputs. This dataset comprises five representative categories related to real-world visual information: (1) Violence, (2) Financial, (3) Privacy, (4) Self-Harm, and (5) Animal. Each category contains 150 queries, resulting in a total of 750 harmful queries. For evaluating CS-DJ, we employed LLaVA-CC3M-Pretrain595K  [^1] https://huggingface.co/datasets/liuhaotian/LLaVA-CC3M-Pretrain-595K to retrieve the contrasting images. To improve retrieval efficiency, we randomly selected a subset of 10,000 images for the retrieval dataset, using a different random seed for each experiment to ensure diversity. All harmful queries used throughout the experiments were derived from the Hades dataset.
Victim Models. In our main experiments, We evaluated the performance of the four most popular closed-source MLLMs: GPT-4o-Mini, GPT-4o, GPT-4V [1], and Gemini-1.5-Flash [30]. The specific versions are as follows: GPT-4o: GPT-4o-2024-08-06, GPT-4o-Mini: GPT-4o-Mini-2024-07-18, GPT-4V: GPT-4-1106-vision-preview, and Gemini-1.5-flash: Gemini-1.5-flash-001  [^2] Since Gemini-1.0-Pro-Vision has been deprecated on July 12, 2024, we used its upgraded version, Gemini-1.5-Flash..
Evaluation Metrics. To assess the effectiveness of our proposed CS-DJ framework, we employ two primary evaluation metrics: Attack Success Rate (ASR) and Ensemble Attack Success Rate (EASR). ASR quantifies the proportion of successful jailbreak attempts by determining whether the model’s response meets predefined conditions for harmful content. This determination is guided by a security discriminator that classifies responses as harmful or benign based on specific criteria. We utilize Beaver-Dam-7b [7], trained on high-quality human feedback data, as the security discriminator, leveraging its ability to detect both the harmfulness and helpfulness of responses to their corresponding queries. This ensures a robust and consistent evaluation of attack success across various scenarios. EASR [40] evaluates the effectiveness of a set of jailbreak templates by measuring the proportion of queries for which at least one template within the subset successfully bypasses the target MLLM’s defenses. While ASR focuses on the success rate of an individual template, EASR provides a broader assessment by considering the combined success of a group of templates, offering deeper insights into the MLLM’s susceptibility to jailbreak attempts.
4.2 Main Experiment
To assess the effectiveness of CS-DJ, we conducted a comparison between CS-DJ and Hades [10], the state-of-the-art closed-source MLLM jailbreak attack. For a fair comparison, we adopted the official implementation of Hades and performed 6 groups of experiments to collect data.
Table 1 presents the jailbreak ASR and EASR results of CS-DJ and Hades across the four widely used closed-source MLLMs. The results across the five categories indicate that the four closed-source MLLMs exhibit stronger defenses against harmful queries in the Animal and Self-Harm categories while showing greater vulnerability in the Financial, Privacy, and Violence categories. Compared to their ASR results, Hades and CS-DJ achieve average improvements of 10.22% and 21.70% in EASR, respectively, indicating that increasing the number of trials can enhance the jailbreak success rate. Additionally, for both ASR and EASR, CS-DJ consistently outperforms Hades across all five categories for each evaluated MLLM. Overall, compared to Hades, CS-DJ achieves maximum improvements of 54.91% in ASR and 66.67% in EASR. Furthermore, CS-DJ attains average success rates of 52.40% for ASR and 74.10% for EASR. These observations highlight CS-DJ’s superior effectiveness as a jailbreak method for bypassing the defenses of closed-source MLLMs. Jailbreaking examples can be found in the appendix.
A detailed analysis of Hades’ results across the four models reveals that GPT-4V outperforms the other three models in both ASR and EASR. In contrast, the results of CS-DJ indicate that Gemini-1.5-Flash achieved the highest performance. These findings suggest that the impact of visual input harmfulness and distraction varies across different models. In a word, the results demonstrate that distracting the model’s attention is a more effective strategy for enhancing jailbreak success rates.
For the subsequent experiments, we selected the most widely used model, GPT-4o, for testing. All experiments were conducted as a single trial. Therefore, only ASR results are presented in the following sections.
4.3 Impact of Query Decomposition
Setting Ani. Fin. Priv. Self-H. Viol. ASR (%)
RQ 0.67 6.00 5.33 0 4.00 3.20
3SQ 4.00 28.66 34.66 4.66 22.00 18.80
6SQ 8.00 45.33 44.00 10.66 41.33 29.86
9SQ 4.66 44.66 44.66 7.33 36.00 27.46
Table 2: GPT-4o’s ASR results of CS-DJ under different query decomposition settings. RQ denotes using the image transformation of raw queries as the visual input for MLLMs. 3SQ, 6SQ, and 9SQ indicate that the raw query is decomposed into 3, 6, and 9 sub-queries, respectively, which are then transformed into sub-images to compose the final visual input.
Here, we aim to explore the impact of query decomposition on MLLM jailbreak performance. To validate its effectiveness, we first use images derived from raw queries (RQ) as visual inputs for MLLMs, serving as a baseline for comparison with the results obtained using query decomposition. We further examine the effect of the number of sub-queries on the jailbreak success rate by decomposing the raw query into 3, 6, and 9 sub-queries, denoted as 3SQ, 6SQ, and 9SQ, respectively. Given that MLLMs typically resize image inputs to fixed dimensions, we aim to minimize distortion from aspect ratio changes during resizing and avoid potential limitations of the MLLM’s image encoder. Therefore, we maintain a fixed column count of 3 when composing the final input image. As a result, the 3SQ, 6SQ, and 9SQ settings increase the number of rows while keeping the number of columns constant.
The quantitative results are presented in Table 2. The results from RQ demonstrate that the MLLM exhibits strong defenses against original jailbreak queries. Even when the raw query is transformed into an image input, it remains highly challenging to bypass the MLLM’s security mechanisms. In contrast, 3SQ outperforms RQ across all five categories, yielding an overall improvement of 15.60%. This indicates that query decomposition effectively enhances the jailbreak success rate for MLLMs. By introducing a distributional shift in the queries, query decomposition creates structural distraction, thereby enabling more effective bypassing of the model’s defenses. Furthermore, increasing the number of decomposed sub-queries to 6 results in an additional 11.06% improvement in the jailbreak success rate. This suggests that a higher number of sub-images transformed from sub-queries can further enhance the distraction effect on the MLLM. However, when the number of sub-queries is increased to 9, a slight decline in the success rate is observed. This decrease may be attributed to two factors: (1) the increased number of sub-queries raises task complexity, resulting in greater demands on the model’s comprehension and its ability to handle complex tasks; (2) limitations in the model’s image encoding capabilities may prevent accurate interpretation of the sub-queries; and (3) over-decomposition may cause the model to recognize the original intent of the query, thereby reducing the effectiveness of the jailbreak attempt.
Overall, the structural distraction introduced by query decomposition can enhance the jailbreak success rate of MLLMs. Furthermore, the number of sub-queries also plays a significant role in influencing the results. Given the limited capabilities of smaller MLLMs and the need to incorporate additional visual subimages, this work focuses on evaluating the performance of the CS-DJ with three sub-queries.
4.4 Impact of Contrasting Visuals
Figure 3: GPT-4o’s evaluation results for varying quantities of visual subimages.
To validate the effectiveness of contrasting visuals, we conduct experiments from two perspectives: (1) the impact of distraction induced by varying the number of subimages, and (2) the effect of inter-subimage distraction with a fixed number of subimages. Since altering the number of subimages directly affects the layout of the final composite image, we maintain a fixed column count of 3 and adjust the number of rows to control the number of subimages. This strategy reduces distortion due to excessive changes in aspect ratio and mitigates potential limitations in the MLLM’s image encoding capabilities. We conduct different experiments with various configurations, involving 0, 3, 6, 9, and 12 Contrasting visual SubImages, corresponding to 3SQ, 3SQ+3CSI, 3SQ+6CSI, 3SQ+9CSI, and 3SQ+12CSI, respectively. Furthermore, to explore the influence of inter-subimage distraction on jailbreak success rates, we conduct experiments using three settings: retrieving 9 most contrasting images, retrieving 9 most similar images, and using a single, most similar image from the dataset for all 9 subimages, denoting 3SQ+9CSI, 3SQ+9SSI, and 3SQ+9SinSI, respectively.
The experimental results on the impact of visual subimage quantity are illustrated in Figure 3. As 3SQ does not contain any visual subimages, its Distraction Distance is 0, resulting in the lowest ASR. Increasing the number of visual subimages leads to a rise in ASR. Specifically, compared to 3SQ, 3SQ+3CSI obtains a 12.8% increase in ASR, demonstrating that contrasting visual subimages can effectively distract the MLLM and bypass its internal defense mechanisms. As the number of visual subimages increases, both the Distraction Distance and the ASR grow correspondingly. However, the rate of ASR improvement diminishes as the number of subimages continues to rise. When the number of subimages reaches 9, further additions yield limited gains in ASR. Given the limited encoding capabilities of the visual encoders in smaller MLLMs, we use 9 visual subimages as the default setting for our experiments.
Table 3 presents the experimental results on the impact of inter-subimage distraction. We evaluate three different subimage selection strategies. The results show that 3SQ+9SinSI has the lowest distraction distance, corresponding to the lowest ASR. In contrast, 3SQ+9CSI achieves an 11.20% improvement in ASR compared to 3SQ+9SSI, along with a further increase in distraction distance. These findings indicate that, for a fixed number of subimages, greater distraction among subimages leads to a higher jailbreak success rate for MLLMs. Additionally, an analysis of the results from Figure 3 and Table 3 reveals that distraction distance accurately reflects changes in ASR trends only when varying the subimage construction strategy along a single dimension, such as the number of subimages or subimage selection strategies. This limitation may stem from differences between the semantic space of our current feature extractor, CLIP, and that of the MLLM. As a result, inter-subimage distraction distance cannot always be accurately measured using the CLIP model. Future work will focus on developing a more comprehensive metric for measuring distraction distance.
Table 4 shows the impact of information complexity of visual subimages. The ASR of 3SQ+9RNI is close to that of 3SQ, but significantly lower than 3SQ+9CSI, indicating that MLLMs exhibit strong recognition capabilities for noise images with minimal informational content. Therefore, only subimages with higher information complexity can effectively distract the model.
In summary, constructing multiple contrasting visual subimages effectively enhances the distraction of MLLMs, leading to improved jailbreak success rates.
Setting ASR (%) Distraction Distance
3SQ+9SinSI 24.66 399.08
3SQ+9SSI 32.53 1378.77
3SQ+9CSI 43.73 1878.83
Table 3: ASR of GPT-4o and Distraction Distance for CS-DJ under different visual subimage selection strategies. 9SinSI indicates using a single, most similar image from the dataset for all 9 subimages. 9CSI and 9SSI denote retrieving 9 most contrasting and similar images to compose the final visual input, respectively.
Setting Ani. Fin. Priv. Self-H. Viol. ASR (%)
3SQ 4.00 28.66 34.66 4.66 22.00 18.80
3SQ+9RNI 5.33 32.66 33.33 4.66 24.00 20.00
3SQ+9CSI 19.33 73.33 56.66 18.00 51.33 43.73
Table 4: GPT-4o’s ASR results of CS-DJ under different settings. 9RNI denotes using 9 random noise images to compose the final visual input.
4.5 Impact of Instruction
Setting Ani. Fin. Priv. Self-H. Viol. ASR (%)
task-guiding 10.00 55.33 39.33 13.33 42.66 32.00
+ role-guiding 15.33 67.33 50.00 13.33 52.00 39.73
+ visual-guiding 19.33 73.33 56.66 18.00 51.33 43.73
Table 5: GPT-4o’s ASR results of CS-DJ under the 3SQ+9CSI configuration with different instruction $P$ settings.
To complement the constructed multi-subimage input, we carefully designed the instruction $P$. To evaluate the impact of different components of $P$, we conducted three experiments: (1) using only the task-guiding component, (2) adding the role-guiding component, and (3) further incorporating the visual-guiding component. The results, presented in Table 5, show that using only the task-guiding component achieved the lowest ASR but still outperformed Hades by 26.49%. This indicates that the visual distraction from the multi-subimage input, combined with the task distraction of handling multiple tasks simultaneously, significantly contributes to improving jailbreak success rates. Adding the role-guiding component further increased the ASR by 7.73%, while incorporating the visual-guiding component led to an additional 4.00% improvement. These findings demonstrate that targeted optimization of $P$ for multi-subimage visual inputs can further disperse the model’s attention, enhancing the success rate of the jailbreak.
5 Conclusion
In this paper, we investigate the visual vulnerabilities in jailbreak attacks on MLLMs. Our findings show that the complexity of visual inputs significantly influences attack success. We introduce the Distraction Hypothesis and propose the CS-DJ framework, which leverages distraction-based strategies to exploit these vulnerabilities in MLLMs. Our approach integrates this insight by two core components: structured distraction through query decomposition and visual-enhanced distraction using contrasting subimages. Extensive experiments conducted on five representative scenarios and four popular closed-source MLLMs demonstrated that CS-DJ outperforms state-of-the-art jailbreak methods, validating the efficacy of distraction-based approaches in bypassing MLLM defenses.